OpenStudy (anonymous):

Help ASAP ! Will FAN & Medal!! If x = 24 squared and y = 5. What is the length of z?

OpenStudy (anonymous):



OpenStudy (anonymous):

Answers : 49cm 7cm 1 cm 29 squared cm

OpenStudy (zzr0ck3r):

well \(y^2+x^2 = z^2\) so \(z = \sqrt{x^2+y^2}=\sqrt{(24^2)^2+5^2}\)

OpenStudy (zzr0ck3r):

none of those are right

OpenStudy (anonymous):

those are ther answer choices

OpenStudy (anonymous):

the*

OpenStudy (zzr0ck3r):

\(x=24^2\)?

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

ahh this is square rood of 24 not 24 squared

OpenStudy (anonymous):

ye same as the question it's the same 24

OpenStudy (zzr0ck3r):

\(z = \sqrt{x^2+y^2}=\sqrt{\sqrt{24}^2+5^2}=\sqrt{24+25}=\sqrt{49}= ?\)

OpenStudy (zzr0ck3r):

you wrote 24 squared in the question, you mean square root of 24

OpenStudy (anonymous):

yes** 7

OpenStudy (zzr0ck3r):

correct